OpenStudy (anonymous):

how do you work this? [(5x+15)/(10x-10)] / [(x^2+6x+9)/(3x^2-3)]

OpenStudy (bahrom7893):

well simplify each fraction separately

OpenStudy (bahrom7893):

will make life much easier

OpenStudy (mayliz):

((5 x+15) (3 x^2-3))/((10 x-10) (x^2+6 x+9))

OpenStudy (anonymous):

I know. I can simplify everything except the 3x^2-3. Can you simplify it?

OpenStudy (bahrom7893):

yea it becomes 3(x-1)(x+1)

OpenStudy (bahrom7893):

well if u factor it

OpenStudy (anonymous):

Oh, got it. Thanks.

OpenStudy (mertsj):

\[\frac{5(x+3)}{10(x-1)}\times\frac{3(x-1)(x+1)}{(x+3)(x+3)}=\frac{3(x+1)}{2(x+3)}\]

OpenStudy (bahrom7893):

Merts is awesome